OpenStudy (anonymous):

what is the Maximum and minimum of the f(x,y,z)=x^2 -2yz in {(x,y,z):z>=0,x^2+y^2+(z-1)^2<=3

OpenStudy (anonymous):

\[|x|_{\max}=|y|_{\max}=|z-1|_{\max}=\sqrt{3}\] zmax=sqrt(3)-1 and z min=0 \[x_{\max}=\sqrt3 ,x_{\min}=-\sqrt3;\]\[y_{\max}=\sqrt3 ,y_{\min}=-\sqrt3\] so f(x,y,z)_{max}=(xmax)^2-2*ymin*zmax =3+2*sqrt(3)*(sqrt(3)-1) =9-2sqrt(3)

OpenStudy (anonymous):

f_{min}=0^2-2*ymax*zmax =-2*sqrt(3)*(sqrt(3)-1) =-6+sqrt(3)

OpenStudy (anonymous):

INTERNAL EXTREMA: Partial Derivatives: fx = 2x fy = -2z fz = -2y Set each to zero and you get -> P(0,0,0) f(0,0,0) = 0 BOUNDARY EXTREMA: Because the region is a circle (center (0,0,1); radius sqrt(3)) your max will be along the boundary. Using common sense you can tell this will happen at: \[x = \sqrt3\] and f(sqrt(3),0,0) = 3 which is a max. f(0,sqrt(3),0) and f(0,0,sqrt(3)) are both -2sqrt(3) corresponding to a minumum. therefore the point found earlier, (0,0,0) is a saddle point.